OpenStudy (anonymous):

z2 – z – 56 how to factor completely this problem?

OpenStudy (anonymous):

(z - 8)(z + 7)

OpenStudy (anonymous):

Which gives you z = 8 and z = -7 as your roots.

OpenStudy (anonymous):

I get the opposite (z-7)(z+8) are they the same?

OpenStudy (anonymous):

no they are not the same. FOIL your answer you would get z^2 + z - 56

OpenStudy (anonymous):

you want two numbers that multiply to -56 and add up to -1 -8 and 7 will do that 8 and -7 will get you a sum of +1 HOWEVER, the ROOTS of the equation, as BubbaMurphy says are 8 and -7 because you have (z - 8)(z + 7) = 0 so you get z - 8 = 0 z + 7 = 0 giving you z = 8 and -7

OpenStudy (anonymous):

oh okey thanks guys u save me

OpenStudy (anonymous):

No problem